OpenStudy (anonymous):

Solve the following equation for x. -2(17x-4)+7(5x-7)-9=2-5

OpenStudy (blurbendy):

-2(17x-4)+7(5x-7)-9=2-5 -34x + 8 + 35x -49 - 9 = - 3 x - 49 + 8 -9 = -3 x = 49 - 8 + 9 - 3 Can you solve from here?

OpenStudy (amoodarya):



OpenStudy (anonymous):

Thank you!

OpenStudy (amoodarya):

solve it ?